OpenStudy (anonymous):

What is the average of four tenths and five thousandths? A. 25002 B. 2502 C. 0.225 D. 0.2025 E. 0.02025

OpenStudy (mathstudent55):

Add them together and divide by 2.

OpenStudy (mathstudent55):

Here's one way of doing it: \( \dfrac{\frac{4}{10} + \frac{5}{1000}}{2} \) Here's another way: \( \dfrac{0.4 + 0.005}{2} \)

OpenStudy (anonymous):

0.2025

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

Thank u

OpenStudy (mathstudent55):

You're welcome.